# Towards Understanding 3D Vision: the Role of Gaussian Curvature

### v1

@@ Abstract
Recent advances in computer vision have predominantly relied on data-driven approaches that leverage deep learning and large-scale datasets. Deep neural networks have achieved remarkable success in tasks such as stereo matching and monocular depth reconstruction. However, these methods lack explicit models of 3D geometry that can be directly analyzed, transferred across modalities, or systematically modified for controlled experimentation. We investigate the role of Gaussian curvature in 3D surface modeling. Besides Gaussian curvature being an invariant quantity under change of observers or coordinate systems, we demonstrate using the Middlebury stereo dataset that it offers: (i) a sparse and compact description of 3D surfaces, (ii) state-of-the-art monocular and stereo methods seem to implicitly consider it, but no explicit module of such use can be extracted, (iii) a form of geometric prior that can inform and improve 3D surface reconstruction, and (iv) a possible use as an unsupervised metric for stereo methods.
@@ 1 Introduction
The world around us is shaped both by natural processes and by human-made structures, which are commonly represented in stereo vision by depth and disparity measurements. These metrics can accurately describe 3D geometry but depend heavily on the observer’s position and viewpoint, limiting their robustness. In contrast, Gaussian curvature is a purely local geometric property of a surface that remains invariant under changes in viewpoint, making it ideal for robust 3D surface analysis.
Our main contribution is on Foundational Vision Understanding by deepening our understanding of 3D scene geometry, paving the way for more interpretable, generalizable, and reliable vision systems. This is an analytical work grounded on data from state-of-the-art (SOTA) DL stereo algorithms. We investigate their benchmarking performance and how this insight about the importance of Gaussian curvature description of the 3D world is implicitly captured by stereo algorithms with indoor scenes on Middlebury dataset. We show that (i) Gaussian curvature is sparsely distributed across natural 3D surfaces; (ii) Low Gaussian curvature can serve as a prior for regularizing or modeling 3D surface geometry; (iii) This prior may be implicitly captured by current SOTA algorithms, but it is not explicit as an independent module and so we cannot use it elsewhere; (iv) Low Gaussian curvature enables the definition of a novel unsupervised evaluation metric for assessing the quality of 3D surface reconstruction algorithms.
@@
Scene Understanding: an analysis of the role of Gaussian curvature in 3D data representations of indoor environments. The identification of quantities that are zero or nearly zero across most of a scene – thus encoding data with a minimal set of active components – forms the foundation of sparse representation.
Explainable AI: an analysis of the geometric constraints learned by SOTA stereo algorithms suggests they rely on low Gaussian curvature priors. However, it is unclear if or where these priors are explicitly used, making it impossible to isolate reusable algorithmic modules from current depth reconstruction methods. A potential immediate application of this explainability study is using it as an unsupervised metric to evaluate stereo algorithms.
1.1 Previous Work
@@
In contrast to previous approaches, which often focus on quantities defined within specific coordinate frames or based on appearance [27, 20], our work seeks a geometric quantity that is invariant to the choice of observer or coordinate system, a property that is especially desirable for 3D surface analysis. To the best of our knowledge, there have been no prior studies applying sparse representation principles to Gaussian curvature in 3D geometry.
@@ 2.2 Methods to estimate GC from data
where the fundamental forms I and II can be obtained from a parametric surface $\mathbf{Z}(u,v)$, its derivatives $\mathbf{Z}_{u}(u,v),\mathbf{Z}_{v}(u,v)$ and second derivatives $\mathbf{Z}_{uu},\mathbf{Z}_{uv},\mathbf{Z}_{vv}$. More details of the formula are in the supplementary material. We construct the parametrized surface from the depth data. More precisely, given a set of $n$ depth measurements $\{d_{i}(x_{i},y_{i})\}_{i=1}^{n}$, we represent the surface as a 3D point cloud $\{(X_{i},Y_{i},d_{i}(x_{i},y_{i}))_{i=1}^{n}\}$ where $X_{i},Y_{i}$ are obtained by inverting the projection, that is, $X_{i}=\frac{(x_{i}-c_{x})}{f_{x}}d_{i}(x_{i},y_{i})$ and $Y_{i}=\frac{(y_{i}-c_{y})}{f_{y}}d_{i}(x_{i},y_{i})$, before assigning a parametrized surface $\mathbf{Z}(u,v)$. Since numerical differentiation tends to amplify noise, especially when applied to noisy depth data, we apply smoothing when necessary. Importantly, we smoothed the 3D point cloud with Gaussian smoothing over each 3D component, $X,Y,Z$, independently. Note that one should not smooth the raw depth map, as the depth values are not uniformly sampled with respect to surface distances, and the smoothing in the image space does not adequately account for the true geometry (see Figures 1 and 2).
@@
In order to evaluate algorithms that estimate GC on shapes we created five 3D synthetic scenes1113D Synthetic Scenes Page: https://github.com/SherlonAlmeida/Stereo-Cameras-Simulation composed of developable surfaces, such as cylinders, boxes, and planes, which have zero GC, and spheres representing a convex surface with constant positive curvature. The simulated environment was created using Unity 3D. We created two physical cameras inspired on Middlebury settings, with Sensor Size (H,W) = (14.8, 22.2)mm, Image Size (H,W) = (2000, 3000)px, Focal Length (mm, px) = (35, 4729.73) and Pixel Size = 0.0074mm. The cameras were YZ-aligned, and X-shifted by 200 millimeters to guarantee the images are rectified. Thus, Baseline = 200mm.
@@ 3 A Sparse Representation of 3D data
We investigate a sparse representation associated with depth data by analyzing the Middlebury stereo dataset [18], which provides accurate ground truth depth maps. The Middlebury dataset provides high-resolution stereo image pairs along with Ground Truth (GT) disparity maps with sub-pixel accuracy. It is used as a well-established benchmark for evaluating the performance of stereo algorithms. The dataset contains $15$ images, each of size approximately $2,000\times 3,000$ pixels, yielding on the order of $9\times 10^{7}=15\times 2,000\times 3,000$ ground truth depth values. For each pixel, we compute the GC (throwing away boundary data) and aggregate the results and normalize them into a histogram, shown in Figure 5.
@@
A natural metric to evaluate the sparsity of the GC distribution in an image or dataset is the Shannon entropy of the distribution given by equation (3). Alternatively, we propose a simpler to compute and interpretable metric, which we call Low Gaussian Curvature (LGC) metric, where a low entropy should map to a high LGC. This metric is defined as the percentage of all computed GC values that lie within a specified range, [W,W][-W,W]. To avoid the impact of high values near surface boundaries, we remove the top 20% of the highest $|K|$ values from the entire analysis. As shown in Figure 5, most of the GC values in the Middlebury dataset are concentrated in the range of $W=1{,}000\,\mathrm{m}^{-2}$. After applying smoothing with $\sigma=2\,m$ to the data, the LGC increases to 96.3% of values that fall within $1{,}000;1{,}000]\,\mathrm{m}^{-2}$, while the maximum absolute curvature reaches $|K|_{\text{max}}=32{,}370.4\,\mathrm{m}^{-2}$.
@@ 4 An analysis on depth reconstruction
We design experiments using a GC-based metric to evaluate the performance of depth-reconstruction algorithms. These experiments are carried out on indoors scenes from both real-world data, using the Middlebury dataset (subsection 4.1), and controlled synthetic environments from our created 3D synthetic scenes (subsection 4.2).
@@ 4.1 Middlebury Dataset
The dataset includes a training partition, where GT data is publicly available, and a test partition, where only some calibration information and the left-right image pairs are provided. In both partitions, algorithm rankings are available; however, only in the training set can the disparity map results of submitted methods be downloaded for further analysis.
@@
We obtained the full-resolution results of SOTA techniques [^2] There are other relevant methods reported in the literature; however, our selection was based on top-ranked methods on the Middlebury benchmark and the availability of source code., organized into two groups: Group A) FoundationStereo [22], LG-Stereo [24], DEFOM-Stereo [12], and MonoStereo [3], which represent the new generation of stereo approaches that combine stereo matching with monocular features; and Group B) RAFT-Stereo [15], Selective-IGEV [21], DLNR [25], and CREStereo [14], corresponding to previous SOTA techniques.
We evaluated the GC of the 15 training images for both the GT and all previously mentioned approaches. Middlebury results are provided as disparity maps, so we computed $Depth=\frac{f\cdot b}{d+doffs}$, where $f$ is the focal length in pixels, $b$ is the baseline in meters, and $d+doffs$ is the disparity value in pixels corrected by the center of projection offset.
@@
The experiment on Figure 6 compares the disparity AvgError performance and the median $K$ values of GC for all techniques. Note that the new SOTA methods (Group A) are the best-performing approaches in the Middlebury ranking and tend to estimate lower curvature magnitudes compared to the Group B methods.
@@
Figure 6: Middlebury Dataset - Curvature x AvgError Analysis: Top: shows the overall position of each approach in the Middlebury ranking, along with its average disparity error (AvgError). Bottom: presents the median GC for all techniques and the GT.
@@
A more detailed experiment can be seen on Figure 7 and Table 1, where we analyze the GC distribution for each technique. Note that FoundationStereo has 81.3% of its GC between $1{,}000,1{,}000]m^{-2}$ $K$ values. Additionally, the same trend of minimizing GC across the 15 Middlebury training set was observed in all Group A techniques [^3] We invite the reader to refer to the Supplementary Material for a more comprehensive study.. Despite achieving low AvgError, techniques from Group B exhibit also lower LGC, with a tendency toward higher $|K|$ values compared to those observed in Group A.
@@
Figure 7: Curvature Distribution: each plot presents a normalized histogram of the GC distribution for all 15 training images from the Middlebury dataset. We discarded the highest 20% of $|K|$ values, and plotted the remaining $K$ values within $4,000,4,000]m^{-2}$ in 50 bins uniformly distributed.
Technique LGC AvgErr $\downarrow$ RMS $\downarrow$ Bad 0.5 $\downarrow$ Bad 2.0 $\downarrow$
FoundationStereo [22] 81.3 $0.33^{1}$ $2.86^{2}$ $9.46^{1}$ $0.79^{1}$
LG-Stereo [24] 78.1 $0.40^{3}$ $2.94^{3}$ $12.9^{2}$ $1.04^{3}$
DEFOM-Stereo [12] 69.2 $0.61^{5}$ $3.66^{5}$ $19.4^{4}$ $2.26^{5}$
MonoStereo [3] 64.6 $0.34^{2}$ $2.46^{1}$ $13.6^{3}$ $0.94^{2}$
RAFT-Stereo [15] 29.6 $1.04^{8}$ $5.25^{8}$ $28.6^{8}$ $5.25^{8}$
DLNR [25] 36.3 $0.67^{6}$ $3.90^{6}$ $21.5^{6}$ $2.92^{6}$
CREStereo [14] 25.8 $0.94^{7}$ $5.21^{7}$ $26.2^{7}$ $4.01^{7}$
Selective-IGEV [21] 14.0 $0.56^{4}$ $3.57^{4}$ $20.2^{5}$ $2.18^{4}$
Ground Truth 76.3
@@
Table 1: This table presents the Middlebury benchmark ranking for the 15 training images, with techniques listed in descending LGC order. Superscripts indicate each method’s rank among the compared techniques for the metrics AvgError, RMS, Bad 0.5, and Bad 2.0. Darker cell shading highlights better performance. Notably, top-performing methods (Group A) generally exhibit higher LGC (i.e., lower GC).
@@
Considering that humans easily perceive the shape of flat objects (e.g. walls, ground, doors, etc), in the following experiment we present a visual analysis of curvature, using the “Piano” data from Middlebury dataset. We compared the $K$ values of GC from the Middlebury’s provided GT and the FoundationStereo reconstruction, which is the best (lowest AvgError) on the Middlebury ranking for this data. We show the left image, and depth for GT and FoundationStereo in Figure 8, emphasizing the piano bench and sheet music, that we known are developable surfaces with $K=0$. Despite good depth estimation, Middlebury’s GT provides a binary mask with measurement problems for each disparity map. Besides filtering the highest $|K|$ values of the GT, previously we removed these NaN values from the GC analysis for GT. An example of these problematic data can be seen as black positions on the GT depth in Figure 8.
@@
Figure 10 presents a quantitative analysis of GC and average depth error across the five scenes in the 3D synthetic scenes. The left-most bar in each graph represents the curvature of the GT, while the other bars show the curvature estimated by stereo and monocular methods. The right-most plot for each scene shows the average depth error (AvgErr) in centimeters. Notably, FoundationStereo, which achieves the lowest average error – less than one centimeter in all scenes – also exhibits the lowest GC. See the depth and curvature results for each one of them in Figure 11.
Figure 10: This figure presents an quantitative analysis on the average $K$ values of GC (left) and the depth (cm) AvgError performance (right). Observe that stereo approaches with the lowest GC results on the lowest AvgErr. We do not measure AvgErr for monocular approaches since they usually provide relative depth.

### v2

@@ Abstract
Recent advances in computer vision have predominantly relied on data-driven approaches that leverage deep learning and large-scale datasets. Deep neural networks have achieved remarkable success in tasks such as stereo matching and monocular depth reconstruction. However, these methods lack explicit models of 3D geometry that can be directly analyzed, transferred across modalities, or systematically modified for controlled experimentation. We investigate the role of Gaussian curvature in 3D surface modeling. Besides Gaussian curvature being an invariant quantity under change of observers or coordinate systems, we demonstrate using the Middlebury stereo dataset that it offers a sparse and compact description of 3D surfaces. Furthermore, we show a strong correlation between the performance rank of top state-of-the-art stereo and monocular methods and the low total absolute Gaussian curvature. We propose that this property can serve as a geometric prior to improve future 3D reconstruction algorithms.
@@ 1 Introduction
The world around us is shaped both by natural processes and by human-made structures, which are commonly represented in stereo-vision by depth and disparity measurements. These metrics can accurately describe 3D geometry but depend heavily on the observer’s position and viewpoint, limiting their robustness. In contrast, Gaussian curvature is a purely local geometric property of a surface that remains invariant under changes in viewpoint, making it ideal for robust 3D surface analysis.
Our main contribution is on Foundational Vision Understanding by deepening our understanding of 3D scene geometry, paving the way for more interpretable, generalizable, and reliable vision systems. This is an analytical work grounded on data from state-of-the-art (SOTA) Deep Learning (DL) stereo algorithms. We investigate their benchmarking performance and how this insight about the importance of Gaussian curvature description of the 3D world is implicitly captured by stereo algorithms with indoor scenes on Middlebury dataset. We show that (i) Gaussian curvature is sparsely distributed across natural 3D surfaces; (ii) Low Gaussian curvature magnitude can serve as a prior for regularizing or modeling 3D surface geometry; (iii) This prior may be implicitly captured by current SOTA algorithms, but it is not explicit as an independent module and so we cannot use it elsewhere; (iv) Low Gaussian curvature magnitude enables the definition of a novel unsupervised evaluation metric for assessing the quality of 3D surface reconstruction algorithms.
@@
Scene Understanding: The identification of quantities that are zero or nearly zero across most of a scene – thus encoding data with a minimal set of active components – forms the foundation of sparse representation. We show in Section 4 that the Gaussian curvature magnitude fulfills this role for 3D data representations of indoor environments.
Explainable AI: In Section 5 we empirically verify that the SOTA stereo algorithms seem to apply low Gaussian curvature magnitude priors. However, it is unclear whether or where these priors are explicitly used, making it impossible to isolate reusable algorithmic modules, in analogy to modules that perform feature extraction, from current depth reconstruction methods.
A potential immediate application of this explainability study is to use low Gaussian curvature magnitude as an unsupervised metric to regularize models and/or evaluate 3D reconstruction algorithms.
2 Previous Work
@@
In contrast to previous approaches to 3D surface understanding, which often focus on quantities defined within specific coordinate frames or based on appearance [32, 25], our work seeks a geometric quantity that is invariant to the choice of observer or coordinate system, a property that is especially desirable for 3D surface analysis. To the best of our knowledge, there have been no prior studies applying sparse representation principles to Gaussian curvature in 3D geometry.
@@ 3.2 Methods to estimate GC from data
where the fundamental forms I and II can be obtained from a parametric surface $\mathbf{Z}(u,v)$, its derivatives $\mathbf{Z}_{u}(u,v),\mathbf{Z}_{v}(u,v)$ and second derivatives $\mathbf{Z}_{uu},\mathbf{Z}_{uv},\mathbf{Z}_{vv}$. More details of the formula are in the supplementary material.
We construct the parametrized surface from the depth data. More precisely, given a set of $n$ depth measurements $\{d_{i}(x_{i},y_{i})\}_{i=1}^{n}$, we represent the surface as a 3D point cloud $\{(X_{i},Y_{i},d_{i}(x_{i},y_{i}))_{i=1}^{n}\}$ where $X_{i},Y_{i}$ are obtained by inverting the projection, that is, $X_{i}=\frac{(x_{i}-c_{x})}{f_{x}}d_{i}(x_{i},y_{i})$ and $Y_{i}=\frac{(y_{i}-c_{y})}{f_{y}}d_{i}(x_{i},y_{i})$, before assigning a parametrized surface $\mathbf{Z}(u,v)$. Since numerical differentiation tends to amplify noise, especially when applied to noisy depth data, we apply smoothing when necessary. Importantly, we smoothed the 3D point cloud with Gaussian smoothing over each 3D component, $X,Y,Z$, independently. Note that one should not smooth the raw depth map, as the depth values are not uniformly sampled with respect to surface distances, and the smoothing in the image space does not adequately account for the true geometry (see Figures 1 and 2).
@@
In order to evaluate algorithms that estimate GC on shapes we created five 3D synthetic scenes1113D Synthetic Scenes Page: https://github.com/SherlonAlmeida/Stereo-Cameras-Simulation composed of developable surfaces (zero GC), such as cylinders, boxes, and planes, and spheres representing a convex surface with constant positive curvature. Despite the simplicity of the scenes, it provides complete control over stereo image acquisition and precise ground truth for GC analysis.
The simulated environment was created using Unity 3D. We created two physical cameras inspired on Middlebury settings, with Sensor Size (H,W) = (14.8, 22.2)mm, Image Size (H,W) = (2000, 3000)px, Focal Length (mm, px) = (35, 4729.73) and Pixel Size = 0.0074mm. The cameras were YZ-aligned, and X-shifted by 200 millimeters to guarantee the images are rectified. Thus, Baseline = 200mm.
@@ 4 A Sparse Representation of 3D data
We investigate a sparse representation associated with depth data by analyzing the Middlebury stereo dataset [22], which provides accurate ground truth (GT) depth maps.
The Middlebury dataset includes a training partition, where GT data is publicly available. The training dataset contains $15$ images, each of size approximately $2,000\times 3,000$ pixels, yielding on the order of $9\times 10^{7}=15\times 2,000\times 3,000$ ground truth depth values.
Middlebury results are provided as disparity maps, so we evaluated the GC of the 15 training images as follows: we computed $Depth=\frac{f\cdot b}{d+doffs}$, where $f$ is the focal length in pixels, $b$ is the baseline in meters, and $d+doffs$ is the disparity value in pixels corrected by the center of projection offset. For each pixel, we compute the GC (throwing away boundary data) and aggregate the results and normalize them into a histogram, shown in Figure 5.
@@
A natural metric to evaluate the sparsity of the GC distribution in an image or dataset is the Shannon entropy of $h(K)$. Alternatively, we propose a simpler to compute and interpretable metric, which we call Low Gaussian Curvature (LGC) metric, where a low entropy should map to a high LGC. This metric is defined as the percentage of all computed GC values that lie within a specified range, [W,W][-W,W]. To avoid the impact of high values near surface boundaries (depth discontinuities), we remove the top 20% of the highest $|K|$ values from the entire analysis. As shown in Figure 5, 76.3% of the GC values in the Middlebury dataset are concentrated in the range of $W=1{,}000\,\mathrm{m}^{-2}$. After applying smoothing with $\sigma=2\,m$ to the data, the LGC increases to 96.3% of values that fall within $1{,}000;1{,}000]\,\mathrm{m}^{-2}$, while the maximum absolute curvature reaches $|K|_{\text{max}}=32{,}370.4\,\mathrm{m}^{-2}$.
@@ 5 An analysis on depth reconstruction
We design experiments to evaluate to what extent depth-reconstruction algorithms incorporate a sparse representation for the magnitude of the GC. These experiments are carried out on indoors scenes from both real-world data, using the Middlebury dataset and controlled synthetic environments from our created 3D synthetic scenes.
@@ 5.1 Middlebury Dataset
The dataset includes a training partition, where GT data is publicly available, and a test partition, where only some calibration information and the left-right image pairs are provided. In both partitions, algorithm rankings are available; however, only in the training set can the disparity map results of submitted methods be downloaded for further analysis. We obtained the full-resolution results of SOTA techniques [^2] There are other relevant methods reported in the literature; however, our selection was based on top-ranked methods on the Middlebury, KITTI, and ETH3D benchmarks and the availability of source code., organized into two groups: Group A) FoundationStereo [27], LG-Stereo [^3] Papers and code for LG-Stereo, BLMT-Stereo, and DepthFocus are not available yet., DEFOM-Stereo [14], MonoStereo [4], MonSter++ [3], and AIO-Stereo [31] which represent the new generation of stereo approaches that combine stereo matching with monocular features; and Group B) RAFT-Stereo [17], CREStereo [16], DLNR [29], Selective-IGEV [26], and S2M2 [20], corresponding to previous SOTA techniques. We also evaluated the results of BLMT-Stereofootnotemark: and DepthFocusfootnotemark: , two new approaches currently under review for CVPR 2026. Although the papers and source codes were not available at the time of our submission, we were able to analyze their results using the GC evaluation.
@@
The experiment on Figure 6 compares the disparity AvgError performance and the median $|K|$ values of GC for all techniques. Note that the new SOTA methods (Group A) are the best-performing approaches in the Middlebury ranking and tend to estimate lower curvature magnitudes compared to the Group B methods. Table 1 presents a summary of the evaluation metrics.
Metric Definition
Average Absolute Error
$\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}|d_{ij}-\hat{d}_{ij}|$
Root Mean Squared Error
$\sqrt{\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}(d_{ij}-\hat{d}_{ij})^{2}}$
Bad-N
(% of pixels with error >N>N)
$100\times\frac{\big|\{\,(i,j)\;|\;|d_{ij}-\hat{d}_{ij}|>N\,\}\big|}{\big|\{\,(i,j)\;\text{ valid pixels }\}\big|}$
Average Absolute
Gaussian Curvature (Avg|K||K|)
$\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}|K_{ij}|$
Normal Error
$\frac{1}{|\Omega|}\sum_{(i,j)\in\Omega}\left(1-\frac{\mathbf{n}_{ij}^{\text{gt}}\cdot\mathbf{n}_{ij}^{\text{pred}}}{\lVert\mathbf{n}_{ij}^{\text{gt}}\rVert\,\lVert\mathbf{n}_{ij}^{\text{pred}}\rVert}\right)$
where $\Omega$ is the set of valid pixels.
Low Gaussian
Curvature (LGC)
$100\times\frac{\left|\{(i,j)\mid-W\leq K_{ij}\leq W\}\right|}{\left|\{\,(i,j)\;\text{valid pixels}\,\}\right|}$
Table 1: Summary of evaluation metrics adopted in this work, covering disparity and depth errors, curvature measurements, and surface-normal consistency.
@@
Figure 6: Middlebury Dataset - Curvature x AvgError Analysis: Top: shows the overall position of each approach in the Middlebury ranking, along with its average disparity error (AvgError). Bottom: presents the median of absolute GC for all techniques and the GT.
@@
A more detailed experiment can be seen on Figure 7 and Table 2, where we analyze the GC distribution for each technique. Note that FoundationStereo has 81.3% of its GC between $1{,}000,1{,}000]m^{-2}$ $K$ values. Additionally, the same trend of minimizing GC across the 15 Middlebury training set was observed in the Top 5 best techniques, all of them from Group A [^4] We invite the reader to refer to the Supplementary Material for a more comprehensive study.. Despite achieving low AvgError, techniques from Group B exhibit also lower LGC, with a tendency toward higher $|K|$ values compared to those observed in the Top 5 Group A techniques.
@@
Figure 7: Curvature Distribution: each plot presents a normalized histogram of the GC distribution for all 15 training images from the Middlebury dataset. We discarded the highest 20% of $|K|$ values to avoid depth discontinuities, and plotted the remaining $K$ values within $4,000,4,000]m^{-2}$ in 50 bins uniformly distributed.
   Technique    LGC    AvgError $\downarrow$    RMS $\downarrow$    Bad 2.0 $\downarrow$    Bad 4.0 $\downarrow$
   FoundationStereo [27]    81.3    $0.33^{2}$    $2.86^{4}$    $0.79^{2}$    $0.40^{3}$
   LG-Stereo    78.1    $0.40^{5}$    $2.94^{5}$    $1.04^{5}$    $0.44^{4}$
   DEFOM-Stereo [14]    69.2    $0.61^{9}$    $3.66^{8}$    $2.26^{10}$    $1.19^{10}$
   MonoStereo [4]    64.6    $0.34^{3}$    $2.46^{2}$    $0.94^{3}$    $0.35^{2}$
   Monster++ [3]    56.5    $0.39^{4}$    $2.54^{3}$    $1.17^{6}$    $0.47^{5}$
   BLMT-Stereo    43.1    $0.14^{1}$    $1.33^{1}$    $17^{1}    0.0$$09^{1}$
   DepthFocus    39.9    $0.64^{10}$    $4.27^{11}$    $1.84^{7}$    $1.08^{9}$
   DLNR [29]    36.3    $0.67^{11}$    $3.90^{10}$    $2.92^{11}$    $1.41^{11}$
   RAFT-Stereo [17]    29.6    $1.04^{13}$    $5.25^{13}$    $5.25^{13}$    $2.89^{13}$
   CREStereo [16]    25.8    $0.94^{12}$    $5.21^{12}$    $4.01^{12}$    $2.04^{12}$
   S2M2 [20]    17.5    $0.52^{6}$    $3.73^{9}$    $1.00^{4}$    $0.59^{6}$
   AIO-Stereo [31]    15.3    $0.54^{7}$    $3.50^{6}$    $1.97^{8}$    $0.82^{7}$
   Selective-IGEV [26]    14.0    $0.56^{8}$    $3.57^{7}$    $2.18^{9}$    $0.92^{8}$
   Ground Truth    76.3    —    —    —    —
@@
Table 2: This table presents the Middlebury benchmark ranking for the 15 training images, with techniques listed in descending LGC order. Superscripts indicate each method’s rank among the compared techniques for the metrics AvgError, RMS, Bad 2.0, and Bad 4.0. Darker cell shading highlights better performance, indicating the technique is among the Top 1, Top 3, or Top 5 best approaches. Notably, top-performing methods (Group A) generally exhibit higher LGC (i.e., lower GC).
@@
Considering that humans easily perceive the shape of flat objects (e.g. walls, ground, doors, etc), in the following experiment we present a visual analysis of curvature, using the “Piano” data from Middlebury dataset. We compared the $K$ values of GC from the Middlebury’s provided GT and the FoundationStereo reconstruction, which is the best (lowest AvgError) on the Middlebury ranking for this data with public code. We show the left image, and depth for GT and FoundationStereo in Figure 8, emphasizing the piano bench and sheet music, that we known are developable surfaces with $K=0$. Despite good depth estimation, Middlebury’s GT provides a binary mask with measurement problems for each disparity map. Besides filtering the highest $|K|$ values of the GT, previously we removed these NaN values from the GC analysis for GT. An example of these problematic data can be seen as black positions on the GT depth in Figure 8.
@@
In Table 3, we extend our analysis to the normal error between the reconstructed point clouds from each technique and the GT across all 15 Middlebury training images. Notably, the top-performing (Top 5) methods in terms of Normal Average Error (NormAvg) belong to Group A. Moreover, Selective-IGEV showed poor alignment of surface normals with the expected GT across all 15 images.
In Figures 10 and 11, we qualitatively present the normal reconstructions for the GT, BLMT-Stereo, Foundation-Stereo, and Selective-IGEV on the Piano and Playtable data, respectively. BLMT-Stereo and Foundation-Stereo produce depth maps with more coherent surface normals compared to Selective-IGEV. While BLMT-Stereo aligns normals more accurately in thin structures and near depth discontinuities, Foundation-Stereo achieves better consistency in regions with low absolute Gaussian curvature. In both cases, Selective-IGEV struggles, producing surfaces with high absolute GC and poorer normal alignment with the GT.
Technique NormAvg $\downarrow$ Adirondack ArtL Jadeplant Motorcycle MotorcycleE Piano PianoL Pipes Playroom Playtable PlaytableP Recycle Shelves Teddy Vintage
BLMT-Stereo 6.44 $2.09^{1}$ $18.08^{4}$ $4.90^{3}$ $6.31^{1}$ $6.30^{1}$ $2.93^{1}$ $3.01^{1}$ $8.36^{1}$ $3.81^{1}$ $2.60^{1}$ $2.98^{1}$ $1.86^{1}$ $4.40^{1}$ $25.92^{2}$ $3.08^{1}$
FoundationStereo 6.75 $2.24^{2}$ $15.92^{1}$ $4.84^{2}$ $7.88^{4}$ $7.84^{3}$ $3.70^{3}$ $3.97^{3}$ $9.04^{3}$ $4.91^{3}$ $3.09^{3}$ $3.52^{4}$ $2.74^{5}$ $5.03^{3}$ $23.23^{1}$ $3.37^{3}$
LG-Stereo 6.78 $2.09^{1}$ $17.97^{3}$ $4.91^{4}$ $7.00^{2}$ $7.05^{2}$ $3.36^{2}$ $3.68^{2}$ $8.46^{2}$ $4.81^{2}$ $2.92^{2}$ $3.24^{2}$ $2.05^{2}$ $4.90^{2}$ $25.98^{3}$ $3.24^{2}$
MonoStereo 7.29 $2.32^{3}$ $17.75^{2}$ $4.60^{1}$ $7.86^{3}$ $8.32^{4}$ $3.88^{4}$ $4.63^{4}$ $9.92^{4}$ $5.24^{4}$ $14^{4} 3.3$$36^{3}$ $2.36^{3}$ $6.33^{4}$ $26.02^{4}$ $3.64^{4}$
Monster++ 8.13 $2.70^{4}$ $18.26^{5}$ $5.05^{5}$ 8.62 $8.38^{5}$ 4.81 $5.86^{5}$ 11.43 $6.04^{5}$ 4.05 $4.00^{5}$ 2.76 9.57 $26.26^{5}$ 4.19
DEFOM-Stereo 8.18 $2.77^{5}$ 19.18 6.82 $8.56^{5}$ 8.70 4.71 6.18 $11.31^{5}$ 6.53 $3.90^{5}$ 4.15 $2.68^{4}$ $7.16^{5}$ 26.43 $3.66^{5}$
DepthFocus 8.93 4.24 19.38 10.49 8.79 8.83 $4.61^{5}$ 7.85 12.05 6.97 4.01 4.17 3.43 7.35 27.34 4.39
DLNR 9.05 3.50 19.76 7.20 9.40 9.50 6.22 8.88 11.60 6.92 4.20 4.47 3.36 8.68 26.65 5.42
S2M2 9.82 5.51 19.50 6.40 10.63 10.59 5.99 6.07 15.46 8.02 5.82 5.94 4.58 8.54 28.00 6.20
RAFT-Stereo 11.11 4.94 21.57 9.57 11.44 11.55 7.77 11.84 14.67 8.45 5.90 6.16 4.38 12.06 27.08 9.33
CREStereo 11.61 5.69 20.46 10.30 12.55 12.52 9.73 12.94 15.87 9.78 7.80 7.84 5.19 10.17 26.95 6.43
AIO-Stereo 16.65 10.03 22.05 12.28 15.49 15.55 14.60 16.45 23.45 13.21 14.20 14.81 10.46 17.65 28.48 20.99
Selective-IGEV 17.90 10.80 22.62 13.61 17.23 17.28 16.88 18.70 24.09 14.31 15.87 16.44 11.57 18.15 28.98 22.00
DepthPro 19.86 10.74 42.77 31.17 23.37 23.37 11.89 11.89 31.91 18.42 8.87 9.17 9.62 14.88 42.35 7.46
Table 3: This table presents techniques listed in ascending Normal Average Error (NormAvg) order for the 15 Middlebury training images. Superscripts indicate the lowest 1-5th Normal Error (NormalsErr) per technique for each one of the 15 Middlebury images. Darker cell shading highlights lower NormalsErr, indicating the technique is among the Top 1, Top 3, or Top 5 approaches with the lowest NormalsErr. Notably, the top-performing methods are from the Group A, and predict surface normals that are closely aligned with the ground truth.
@@
Figure 10: Normals Analysis: Qualitative comparison on normals reconstruction for Piano data.
Figure 11: Normals Analysis: Qualitative comparison on normals reconstruction for Playtable data.
Figure 12: Quantitative vs. Qualitative comparison: We present the average disparity error (AvgError), average absolute Gaussian curvature (Avg|K||K|), and normal error (NormalsErr) for the Foundation-Stereo and Selective-IGEV techniques on the Adirondack data. We also include a zoomed-in region to illustrate the pointwise consistency of Gaussian curvature and surface normals. Note that despite its low AvgError, the reconstruction produced by Selective-IGEV lacks smoothness.
In Figure 12, we also present additional qualitative and quantitative results for the Adirondack data. Although both Foundation-Stereo and Selective-IGEV achieve accurate depth reconstruction (low AvgError), their Gaussian curvature and normal consistency differ substantially. Foundation-Stereo exhibits lower Avg|K||K| and lower NormalsErr, which may account for its overall better performance.
@@
Figure 13 presents a quantitative analysis of GC and average depth error across the five scenes in the 3D synthetic scenes. The left-most bar in each graph represents the curvature of the GT, while the other bars show the curvature estimated by stereo and monocular methods. The right-most plot for each scene shows the average depth error (AvgError) in centimeters. Notably, FoundationStereo, which achieves the lowest average error – less than one centimeter in all scenes – also exhibits the lowest GC. See the depth and curvature results for each one of them in Figure 14.
Figure 13: This figure presents an quantitative analysis on the average $|K|$ values of GC (left) and the depth (cm) AvgError performance (right). Notice that stereo approaches with the lowest GC values achieve the lowest AvgError. We do not measure AvgError for monocular approaches since they usually provide relative depth.